# The excess of molecular hydrogen in chemical networks without oxygen

### v1

@@ 2.2 Chemical networks
In order to quantify the excess of $\mathrm{H}_{2}$ in reduced networks, we run the model grid calculations for chemical networks with different sets of included elements. The full chemical network within Cloudy consists of 30 elements, their various ionization stages as well as their molecules. This is the reference model and we refer to it as Full. For the reduced chemical networks, we switch off all elements beside a small subset, always including H and He, by adding the element off [element name] commands to the Cloudy input file. Most critically, this does not affect the dust grain composition, which is set independently from the gas phase abundances. For solar metallicity, the formation on the surfaces of dust grains is the dominant formation process of $\mathrm{H}_{2}$, and therefore typically included in the reduced chemical networks mentioned in section 1. By excluding the individual elements from the gas-phase chemistry while keeping the grain properties unchanged, we reproduce the general processes included in reduced networks.
Figure 1: This figure illustrates the two definitions of the $\mathrm{H}_{2}$ excess, $\mathcal{H}$ (top row) and $\mathcal{E}$ (bottom row) for different gas temperatures (columns). The lines in the top row indicate the molecular hydrogen fractions, $\log f_{\mathrm{H2}}\equiv\log(2n_{\mathrm{H2}}/n_{\mathrm{H}})$, for total hydrogen densities, $n_{\mathrm{H}}$, for the reduced (H,He; dashed lines) and the full chemical network (H,He,M; solid lines). The thin horizontal line is at $\log f_{\mathrm{H2}}=-0.5$, and serves as a marker for the transition between atomic and molecular gas. The $\mathrm{H}_{2}$ excess is quantified as $\mathcal{H}$, the difference in the densities at which this transition occurs between the two networks (red horizontal bar) and $\mathcal{E}$, the difference in the $\mathrm{H}_{2}$ fraction at the density of the H i-$\mathrm{H}_{2}$ transition in the reduced HHe network (indicated as blue vertical bar, also in the bottom row). The bottom panels shows $f_{\mathrm{H2}}$ from the reduced networks HHe (black dashed line) and HHeO (grey dotted line) over $f_{\mathrm{H2}}$ from the full network. The $\mathrm{H}_{2}$ excess is drastically reduced in the reduced network HHeO that includes oxygen.
@@ 3 Result
Figure 4: The $\mathrm{H}_{2}$ excess parameter, $\mathcal{H}$, as defined in equation (1) for the model grid described in section 2.1 between the full chemical network and the chemical network that only includes hydrogen and helium species, HHe. $\mathcal{H}$ is evaluated for different gas temperatures (top row: $\log T\,[\mathrm{K}]=2$, middle row: $\log T\,[\mathrm{K}]=2.5$, bottom row: $\log T\,[\mathrm{K}]=3$), gas metallicities (columns from left to right: $\log Z/\mathrm{Z}_{\odot}=-4,-3,-1,0$), and shielding column densities ($\log N_{\mathrm{sh}}$, x-axis). Each panel shows $\mathcal{H}(N_{\mathrm{sh}})$ for different radiation field strengths, from $\log I_{\mathrm{ISRF}}=-2$ to 11 (from dark to light line colors). Missing lines or line fragments indicate that the H i-$\mathrm{H}_{2}$ transition occurs outside the density range of our model grid (see Fig. 3). Individual grid points are missing where Cloudy does not converge.
We use the $\mathrm{H}_{2}$ excess parameters, $\mathcal{H}$ (equation 1), $\log n_{\mathrm{H},\mathcal{E}}$ (equation 2), and $\mathcal{E}$ (equation 3) to quantify the $\mathrm{H}_{2}$ excess across different gas temperatures, metallicities, and radiation field strengths. Fig. 3 shows the density of the H i to $\mathrm{H}_{2}$ transition, for gas temperatures of 100 K ($\log T\,[\mathrm{K}]=2$, top row), 316 K ($\log T\,[\mathrm{K}]=2.5$, middle row), and 1000 K ($\log T\,[\mathrm{K}]=3$, bottom row) and gas metallicities between $\log Z/\mathrm{Z}_{\odot}=-4$ and $\log Z/\mathrm{Z}_{\odot}=0$ (from left to right columns). The dashed lines are for the reduced chemical network HHe and therefore represent $\log n_{\mathrm{H},\mathcal{E}}$ (equation 2), while the solid lines are the transition densities for the full chemistry network (Full). Within each panel, the H i to $\mathrm{H}_{2}$ transition shifts to higher densities for stronger radiation fields (line colors from dark, weak radiation field, to bright, strong radiation field). An offset between the transition densities from the reduced (HHe, dashed) and the full (Full, solid) networks is apparent, especially at high metallicities (third and fourth row), but for warm gas (bottom row) persists at low column densities even for metallicities of $\log Z/\mathrm{Z}_{\odot}=-4$ (first column). For a small subset of grid points, the Cloudy logs showed either warnings or errors, and for these grid points the results are set to nan and therefore appear as missing line fragments.
@@ 3.1 Reducing the H2 excess by adding metal elements
In Ploeckinger2025, the $\mathrm{H}_{2}$ fractions from the full network were largely reproduced when adding oxygen to the primoridal chemical network. We repeat this test here by analysing the $\mathrm{H}_{2}$ excess in Cloudy calculations for a reduced chemical network that includes hydrogen, helium and oxygen species (HHeO, see Table 1). The $\mathrm{H}_{2}$ excess parameter $\mathcal{E}$ for HHeO is shown in Fig. 6. Compared to the results from the primordial network HHe (Fig. 5), the $\mathrm{H}_{2}$ excess is drastically reduced. Only for metallicities of $\log Z/\mathrm{Z}_{\odot}\geq-1$ and a temperature of $1000\,\mathrm{K}$ (bottom row, two rightmost panels) a small, negative $\mathrm{H}_{2}$ excess persists which vanishes only when including both carbon and oxygen (HHeCO, not shown).
We demonstrate in Fig. 7 that including carbon instead of oxygen in the chemical network (together with hydrogen and helium) only marginally reduces the $\mathrm{H}_{2}$ excess. This generalizes the conclusions from Ploeckinger2025 that reactions with oxygen species are largely responsible for reducing the $\mathrm{H}_{2}$ fractions in the full chemistry network, relative to the primordial network. The individual reactions between oxygen species and molecular hydrogen in Cloudy are listed in Table 3 and we investigate their individual contributions for both low and high metallicities in the following sections.
@@ 3.2 Low metallicity
where $k_{\mathrm{H2},x}\,[\mathrm{cm^{3}\,s}^{-1}]$ is the rate coefficient of the reaction between $\mathrm{H}_{2}$ and species $x$, and $n_{x}\,[\mathrm{cm}^{-3}]$ is the number density of species $x$. For the oxygen species that react with $\mathrm{H}_{2}$, the rate coefficients for the neutral species, $\mathrm{O}$, $\mathrm{OH}$, and $\mathrm{O}_{2}$, have a very steep temperature dependence, while the rate coefficients for the ionized species, $\mathrm{O}^{+}$, $\mathrm{OH}^{+}$, and $\mathrm{H}_{2}\mathrm{O}^{+}$, are constant (see e.g. table B1 in Glover2010). At constant temperature, the reaction rates, $R_{x}$, are therefore always proportional to $n_{x}$. In turn, at higher temperatures (e.g. $\log T\,[K]=3$), neutral species may contribute more to the destruction of $\mathrm{H}_{2}$ than at lower temperatures (e.g. $\log T\,[K]=2$), if the decrease in $n_{x}$ is counterbalanced by the drastic increase in $k_{\mathrm{H2},x}$.
@@ 4 Summary
We have shown that the increased $\mathrm{H}_{2}$ fraction (here: $\mathrm{H}_{2}$ excess) in primordial chemical networks found in Ploeckinger2025 using the Chimes chemical network is not limited to specific choices for the radiation fields and shielding column densities. We used Cloudy version 25.00 and found systematic differences in the densities of the H i-$\mathrm{H}_{2}$ transition between primordial and full chemical networks, even if both are supposed to represent the same gas-phase metallicity.
@@
In a forthcoming paper, we will present results from cosmological simulations within the Colibre project (Chaikin2025a; Schaye2025colibre) that use Hybrid-Chimes in both HHe or HHeO mode and discuss the $\mathrm{H}_{2}$ excess in simulated galaxies.
Acknowledgements
I would like to thank Alvaro Hacar and Moa Huppenkothen for insightful discussions. This research was funded by the Austrian Science Fund (FWF) through grant-DOI: 10.55776/V982. The computational results have been achieved using the Austrian Scientific Computing (ASC) infrastructure.

### v2

@@ 2.2 Chemical networks
In order to quantify the excess of $\mathrm{H}_{2}$ in reduced networks, we run the model grid calculations for chemical networks with different sets of included elements. The full chemical network within Cloudy consists of 30 elements, their various ionization stages as well as their molecules. This is the reference model and we refer to it as Full. For the reduced chemical networks, we switch off all elements beside a small subset, always including H and He, by adding the element off [element name] commands to the Cloudy input file.
The dust grain composition in each reduced network matches the corresponding full chemical network comparison calculations. For solar metallicity, the formation on the surfaces of dust grains is the dominant formation process of $\mathrm{H}_{2}$, and therefore typically included in the reduced chemical networks mentioned in section 1. By excluding the individual elements from the gas-phase chemistry while keeping the grain properties unchanged, we reproduce the general processes included in reduced networks.
Figure 1: This figure illustrates the two definitions of the $\mathrm{H}_{2}$ excess, $\mathcal{H}$ (top row) and $\mathcal{E}$ (bottom row) for different gas temperatures (columns). The lines in the top row indicate the molecular hydrogen fractions, $\log f_{\mathrm{H2}}\equiv\log(2n_{\mathrm{H2}}/n_{\mathrm{H}})$, for total hydrogen densities, $n_{\mathrm{H}}$, for the reduced (H,He; dashed lines) and the full chemical network (H,He,M; solid lines). The thin horizontal line is at $\log f_{\mathrm{H2}}=-0.5$, and serves as a marker for the transition between atomic and molecular gas. The $\mathrm{H}_{2}$ excess is quantified as $\mathcal{H}$, the difference in the densities at which this transition occurs between the two networks (red horizontal bar) and $\mathcal{E}$, the difference in the $\mathrm{H}_{2}$ fraction at the density of the H i-$\mathrm{H}_{2}$ transition in the reduced HHe network (indicated as blue vertical bar, also in the bottom panels). The black dashed line in the bottom panels shows $f_{\mathrm{H2}}$ from the reduced network HHe (black dashed line in top panel) relative to the $f_{\mathrm{H2}}$ from the full network (grey solid line in top panel). A thin horizontal line is added to the bottom panels at $f_{\mathrm{H2,HHe}}=f_{\mathrm{H2,Full}}$, i.e. no $\mathrm{H}_{2}$ excess, for reference.
@@ 3 Result
Figure 4: The $\mathrm{H}_{2}$ excess parameter, $\mathcal{H}$, as defined in equation (1) for the model grid described in section 2.1 between the full chemical network and the chemical network that only includes hydrogen and helium species, HHe. $\mathcal{H}$ is evaluated for different gas temperatures (top row: $\log T\,[\mathrm{K}]=2$, middle row: $\log T\,[\mathrm{K}]=2.5$, bottom row: $\log T\,[\mathrm{K}]=3$), gas metallicities (columns from left to right: $\log Z/\mathrm{Z}_{\odot}=-4,-3,-1,0$), and shielding column densities ($\log N_{\mathrm{sh}}$, x-axis). Each panel shows $\mathcal{H}(N_{\mathrm{sh}})$ for different radiation field strengths, from $\log I_{\mathrm{ISRF}}=-2$ to 11 (from dark to light line colors). Missing lines or line fragments indicate that the H i-$\mathrm{H}_{2}$ transition occurs outside the density range of our model grid (see Fig. 3).
We use the $\mathrm{H}_{2}$ excess parameters, $\mathcal{H}$ (equation 1), $\log n_{\mathrm{H},\mathcal{E}}$ (equation 2), and $\mathcal{E}$ (equation 3) to quantify the $\mathrm{H}_{2}$ excess across different gas temperatures, metallicities, and radiation field strengths. Fig. 3 shows the density of the H i to $\mathrm{H}_{2}$ transition, for gas temperatures of 100 K ($\log T\,[\mathrm{K}]=2$, top row), 316 K ($\log T\,[\mathrm{K}]=2.5$, middle row), and 1000 K ($\log T\,[\mathrm{K}]=3$, bottom row) and gas metallicities between $\log Z/\mathrm{Z}_{\odot}=-4$ and $\log Z/\mathrm{Z}_{\odot}=0$ (from left to right columns). The dashed lines are for the reduced chemical network HHe and therefore represent $\log n_{\mathrm{H},\mathcal{E}}$ (equation 2), while the solid lines are the transition densities for the full chemistry network (Full). Within each panel, the H i to $\mathrm{H}_{2}$ transition shifts to higher densities for stronger radiation fields (line colors from dark: weak radiation field, to bright: strong radiation field). An offset between the transition densities from the reduced (HHe, dashed) and the full (Full, solid) networks is apparent, especially at high metallicities (third and fourth row), but for warm gas (bottom row) persists at low column densities even for metallicities of $\log Z/\mathrm{Z}_{\odot}=-4$ (first column).
@@ 3.1 Reducing the H2 excess by adding metal elements
In Ploeckinger et al. (2025), the $\mathrm{H}_{2}$ fractions from the full network were largely reproduced when adding oxygen to the primoridal chemical network. We repeat this test here by analysing the $\mathrm{H}_{2}$ excess in Cloudy calculations for a reduced chemical network that includes hydrogen, helium and oxygen species (HHeO, see Table 1). The $\mathrm{H}_{2}$ excess parameter $\mathcal{E}$ for HHeO is shown in Fig. 6. Compared to the results from the primordial network HHe (Fig. 5), the $\mathrm{H}_{2}$ excess is drastically reduced. Only for metallicities of $\log Z/\mathrm{Z}_{\odot}\geq-1$ and a temperature of $1000\,\mathrm{K}$ (bottom row, two rightmost panels) a small, negative $\mathrm{H}_{2}$ excess persists which vanishes only when including both carbon and oxygen (HHeCO, not shown).
We demonstrate in Fig. 7 that including carbon instead of oxygen in the chemical network (together with hydrogen and helium) only marginally reduces the $\mathrm{H}_{2}$ excess. This generalizes the conclusions from Ploeckinger et al. (2025) that reactions with oxygen species are largely responsible for reducing the $\mathrm{H}_{2}$ fractions in the full chemistry network, relative to the primordial network. The individual reactions between oxygen species and molecular hydrogen in Cloudy are listed in Table 3 and we investigate their individual contributions for both low and high metallicities in the following sections.
@@ 3.2 Low metallicity
where $k_{\mathrm{H2},x}\,[\mathrm{cm^{3}\,s}^{-1}]$ is the rate coefficient of the reaction between $\mathrm{H}_{2}$ and species $x$, and $n_{x}\,[\mathrm{cm}^{-3}]$ is the number density of species $x$. For the oxygen species that react with $\mathrm{H}_{2}$, the rate coefficients for the neutral species, $\mathrm{O}$, $\mathrm{OH}$, and $\mathrm{O}_{2}$, have a very steep temperature dependence, while the rate coefficients for the ionized species, $\mathrm{O}^{+}$, $\mathrm{OH}^{+}$, and $\mathrm{H}_{2}\mathrm{O}^{+}$, are constant (see e.g. table B1 in Glover et al., 2010). At constant temperature, the reaction rates, $R_{x}$, are therefore always proportional to $n_{x}$. In turn, at higher temperatures (e.g. $\log T\,[K]=3$), neutral species may contribute more to the destruction of $\mathrm{H}_{2}$ than at lower temperatures (e.g. $\log T\,[K]=2$), if the decrease in $n_{x}$ is counterbalanced by the drastic increase in $k_{\mathrm{H2},x}$.
@@ 4 Summary
We have shown that the increased $\mathrm{H}_{2}$ fraction (here: $\mathrm{H}_{2}$ excess) in primordial chemical networks found in Ploeckinger et al. (2025) using the Chimes chemical network is not limited to specific choices for the radiation fields and shielding column densities. We used Cloudy version 25.00 and found systematic differences in the densities of the H i-$\mathrm{H}_{2}$ transition between primordial and full chemical networks, even if both are supposed to represent the same gas-phase metallicity.
@@
In a forthcoming paper, we will present results from cosmological simulations within the Colibre project (Chaikin et al., 2025; Schaye et al., 2025) that use Hybrid-Chimes in both HHe or HHeO mode and discuss the $\mathrm{H}_{2}$ excess in simulated galaxies.
Acknowledgments
I would like to thank Alvaro Hacar and Moa Huppenkothen for insightful discussions and the anonymous referee for a constructive report that helped clarify and improve the manuscript. This research was funded by the Austrian Science Fund (FWF) through grant-DOI: 10.55776/V982. The computational results have been achieved using the Austrian Scientific Computing (ASC) infrastructure.